OpenStudy (anonymous):

Alex was finding the average change of his stocks over the following months: -10.29 in June, -8.16 in July, +34.50 in August, +3.12 in September. Estimate the average change in his stocks by rounding first to the nearest whole number and then dividing by the number of months. A. increase of 5 B. decrease of 5 C. increase of 4 D. decrease of 4

OpenStudy (anonymous):

To find the average, just add them all together and divide by the number of things you added: (-10.29) + (-8.16) + (34.50) + (3.12) = -10.29 - 8.16 + 34.50 + 3.12 = 19.17 now divide it by 4, because you added 4 things: (19.17) / (4) = 4.7925 = 4.80 so about an a 5$ increase.

OpenStudy (anonymous):

thank you